please be descriptive so i actually understand it.

A 2.4 kg block slides along a frictionless surface
Answer #1

This is a case of inelastic collision ,where only momentum is conserved always.[not energy]
So,
v1 = 1.3 m/s , v2 = 4.2 m/s
m1 = 2.4 kg, to find , m2= ??
given,
m3 =(m1 +m2) , v3 = 2.3 m/s


So, using, m1v1 +m2v2 = m3v3,
we get, 2.4 x 1.3 + m2 x 4.2 = (2.4+m2) x 2.3
3.12 + m2 x 4.2 = 5.52 + 2.3xm2

1.9 m2 = 2.4
So, m2 = 1.26 Kg --> Answer